• Question 8 Evaluate the limit: lim h 0 (3 + h)2 – 9 h • Question 9 α Evaluate the limit: lim 1 +9 15x + 54 2 - 9
Answer #1

lim (3+h) ²9 hao = h lim (3th 3² hito lim 63+h-3)(3+h+3) hono h lim h 16th) hyo h = lim (6th) huyo =6+ = 6 lim 222_15X+54 - l